# Study of dark counts in optical superconducting transition-edge sensors

### v1

@@ Abstract
Superconducting transition-edge sensors (TESs), known for their high single-photon detection efficiency and low background, are increasingly being used in rare event searches. We present the first comprehensive characterization of optical TES backgrounds, identifying three event types: high-energy, electrical noise, and photon-like. We experimentally verify and simulate the source of high-energy events in optical TESs. We develop methods to isolate photon-like events, the expected signal in dark matter searches, achieving record-low photon-like dark count rates in the eV energy range.
Introduction. In recent years, multilayer dielectric haloscopes have emerged as a promising method for the detection of dark photons (DP), a possible dark matter (DM) candidate [1]. The detection principle of a DP haloscope is as follows: a stack of dielectric bilayers converts dark photons from the local DM halo into Standard Model (SM) photons, then a lens focuses the converted photons onto a photosensor [2, 3]. Since the rest mass of the DM particle is entirely converted into the energy of the SM photon, the choice of the photosensor will determine the DM mass range the haloscope is sensitive to. Two experiments deployed this technology for the first time in 2021: LAMPOST [4], in the USA, and MuDHI [5], in the UAE. MuDHI operated a prototype detector at New York University Abu Dhabi (NYUAD) to search for DP in the mass range around . A commercially available single-photon avalanche diode (SPAD) was used as the photosensor [6]. Although the SPAD could have in principle achieved a higher detection efficiency under optimal conditions (80% at as per the manufacturer’s data sheet), it exhibited a low detection efficiency (2%) when operated near the breakdown voltage, a trade-off needed to reduce the dark count rate (DCR). This DCR was around 100 Hz at temperatures below and when operating a few volts above the breakdown voltage. Such a trade-off between detection efficiency and dark count rate notably reduced the detector’s overall sensitivity.
Quantum sensors, such as superconducting-nanowire single-photon detectors (SNSPDs) used by LAMPOST, offer a significant advantage with their low dark count rate and sensitivity to single-photon energy depositions. A feature of SNSPDs, as well as SPADs, is that they function as an on/off detector [7], only indicating whether the detector has been struck by a particle without measuring its energy. Conversely, superconducting transition-edge sensors (TESs) are micro-calorimeters, whose signal is proportional to the energy deposited by a particle, making them our preferred choice as the photosensor in the dielectric haloscope as they can discriminate events by their energy.
@@
Design and operating principle of TESs. A TES is a thin superconducting film operated at the transition between its superconducting and normal state where it is most sensitive to changes in temperature. The TES working point is stabilized on the transition by negative electrothermal feedback [8] which is achieved by applying a voltage bias. This means that any increase in temperature (due to energy absorption) leads to an increase in resistance, which in turn causes a decrease in current. This drop in current reduces Joule heating, thereby cooling the sensor and bringing it back to the working point. A direct current superconducting quantum interference device (dc-SQUID) connected to the TES acts as a transimpedance amplifier. It converts the current drop resulting from the particle interaction to a measurable voltage drop. This voltage can then be amplified by standard electronic amplifiers and the waveforms can be measured and recorded by an oscilloscope. The voltage drop produced by the dc-SQUID is proportional to the energy deposition in the TES [8].
The two superconducting TESs used in this work were fabricated and operated at the Istituto Nazionale di Ricerca Metrologica (INRiM) in Turin (Italy). They are made of a bilayer of titanium, the superconducting material, and gold, a normal metal to lower the critical temperature, , of the titanium (400 mK) by the proximity effect [9]. Both TESs exhibit a of around and were operated inside an adiabatic demagnetization refrigerator (ADR) with a base temperature of . Since TESs and dc-SQUIDs are sensitive to magnetic fields, including the geomagnetic one, a cryogenic magnetic shield was installed within the ADR around the working area to suppress any magnetic interference that could degrade the sensor’s performance.
Two 20 TESs were deposited on silicon substrates. We refer to the sensors as “TES E” and “TES B.” Most of the data have been acquired from TES E, making it the primary focus of our study. Unless otherwise specified, references to TES data should be assumed to pertain to TES E. In the appendix, we provide details on the TESs fabrication and some information about the characterization and dark count rate of TES B. For both TES E and TES B, the bath temperature was consistently maintained at , with the working point set at 15% of their normal resistance.
Calibration. To calibrate the TES, we coupled it to three different attenuated monochromatic lasers of , , and using an optical fiber. The attenuation allows us to restrict the number of photons arriving onto the TES and avoid signal saturation, focusing on the 1–8 photon region. In Fig. 1, the calibration curve is shown on the right, where we plot the measured signal amplitude from the TES against the known energy of the impinging photon(s). The left side of the figure displays the histograms of counts versus amplitude for each laser wavelength used.
Type of signal events. Signals detected by the TES without any input optics are referred to as “intrinsic detector dark counts” by Ref. [7]. However, in this manuscript we drop the adjective “intrinsic” as we have identified external sources contributing to these counts such that not all of the dark counts come from within the sensor (i.e., they are not intrinsic to it). The observed signals fall into three distinct categories: photon-like, high-energy, and electrical noise events. While the three kinds of pulses are visually distinguishable, we have automated the categorization using the following pipeline developed in Python. After applying a Butterworth filter, we calculate a set of variables for each waveform (e.g., full width at half maximum, maximum amplitude, pulse area, etc.). We employ principal component analysis (PCA) to find the two linear combinations of variables that maximize the variance of the points [10]. K-means clustering is then applied to the resulting 2D phase space to identify clusters of similar events [11]. To ensure the accuracy of this automated sorting, we conduct a manual review of the clusters.
This review process is facilitated by a GUI that allows for the selection of a cluster center, subsequently displaying all waveforms associated with that cluster for quick verification of classification. In some cases, as demonstrated in Fig. 2, it is necessary to reapply the clustering algorithm on a subset of the data to fully separate the three groups.
Photon-like events are the green points in group A in Fig. 2, which was obtained by rerunning the classification algorithm on clusters 8 and 9. These signals are characterized by a sharp rise time and look identical to the ones we see when we shine light from the attenuated laser diode onto the TES.
Signals in clusters 1–7 (as shown in Fig. 2) are characterized by a long rise and decay time, suggesting these events are caused by large energy depositions in the TES substrate, likely coming from primary or secondary cosmic ray particles or naturally occurring radioactive decays. We therefore refer to them as “high-energy events.” This classification is consistent with the observed long rise time, which can be attributed to the slow diffusion of energy from the substrate into the TES.
Electrical noise events (orange cluster B in Fig. 2 obtained by reanalyzing clusters 8 and 9) exhibit voltage spikes oscillating between positive and negative values and are characterized by an amplitude area close to zero. We attribute these events to electromagnetic interference with the readout apparatus. The fact that they are still detectable even when the TES is in a fully superconducting state, a condition where it is not sensitive to energy depositions, indicates that the culprit is not the TES. However, the SQUID remains highly sensitive to external electromagnetic disturbances regardless of the TES’s state.
Experimental runs. Photon-like events are the expected signal for dielectric haloscopes searching for dark matter dark photons. Therefore, it is essential to measure the intrinsic photon-like DCR in the TES (i.e, when the sensor is operated in the dark, with no optical link to ambient and in the absence of radioactive sources) as this forms the background against which the dielectric haloscope operates.
At the onset of our investigation, we expected to see an intrinsic photon-like DCR on the order of Hz, in line with findings by Ref. [12], the only preceding study of a similar nature. There was also a question of whether cosmic rays or natural radioactive decays, interacting with either the TES or its substrate, could account for some of these photon-like events as suggested by the same authors. To explore this possibility, and to gain a deeper understanding of the nature of high-energy events, which have not been extensively studied before, we carried out four specific tests, explained below.
It is important to note that all the experimental runs were conducted in the absence of a light source and without an optical fiber connected to the TES, which was entirely encased in a copper box thermally coupled to the -stage, as can be seen in Fig. 3. In Fig. 4, the lid of the copper box is not present and the substrates on which the TESs are deposited can be seen. Although three fibers were present in the cryostat to allow for the initial calibration procedures, their ends were fixed to the outer side of the -stage within the ADR.
The first was a background test (Run 1) intended to measure the photon-like DCR. Additionally, in the same run, we measured the high-energy DCR, which we hypothesized to be caused by cosmic rays or natural radioactive decays interacting with the TES substrate.
In Run 2 and Run 3, we used two radioactive sources, Th and Na, respectively. These sources were placed against the exterior of the ADR and aligned perpendicularly to the TES substrate. The aim of these tests was to examine the TES’s response to high-energy gamma events. In Run 2, a Th source with an activity measured at approximately was used. For Run 3, we employed a Na source, which had an activity of around . The emission spectrum of Na is characterized by two prominent photoelectric peaks. Notably, one of these peaks corresponds to the simultaneous emission of two back-to-back photons, a phenomenon that occurs due to electron-positron annihilation following a decay. This source was sandwiched between the ADR and a plastic scintillator connected to a PMT, which we call a “saber.” This setup allowed us to measure the double coincidences of the back-to-back gammas, capturing these occurrences with both the saber and the TES.
For the fourth test, to demonstrate that cosmic rays can induce high-energy events in the TES, we implemented a cosmic ray coincidence system, placing three sabers, side-by-side, directly below the ADR. A simultaneous triggering of the TES and all sabers indicated a cosmic ray shower or muon bundle event. The three sabers were operated at the same gain within a 10% uncertainty. The threshold was set at on all three, which corresponds to (see appendix for more details about the sabers’ calibration).
The data acquisition was conducted in two modes: in Run 4A, which lasted about 19 hours, the oscilloscope was configured to trigger on the AND signal of the three sabers, recording the TES signal simultaneously; in Run 4B, spanning around 14 hours, the trigger was set on the TES for signals greater than , while recording all three sabers’ outputs. Run 4A provided the ratio of quadruple coincidences (events registered by both the TES and the three sabers) to triple coincidences (events detected solely by the sabers), while Run 4B allowed us to estimate the ratio of quadruple coincidences to the total number of high-energy events detected by the TES.
GEANT4 simulation. Based on preliminary calculations, it is clear that the frequency of cosmic ray interactions with the TES alone is insufficient to account for the number of high-energy events observed. This suggests that high-energy events observed by the TES originate from energy depositions within the substrate. The substrate for TES E is 6.7 mm2.7 mm in area and in thickness. We developed a GEANT4-based simulation to validate our understanding of the high-energy events seen by the sensor. We ran the simulation in three configurations named Sim 2, Sim 3, and Sim 4 which mirror the experimental setups of Run 2, Run 3, and Run 4, respectively. All simulations used the substrate as the sensitive detector rather than the TES itself.
In Sim 2, we modeled the Th source as in Run 2 and simulated one billion decay events. The simulation assumed the source to be in secular equilibrium. This approach is supported by empirical data obtained from analyzing the thorium source with a high-purity germanium (HPGe) detector at NYUAD, which provided activity level and spectral information consistent with this assumption. The activity from these measurements, around , was also used to convert the number of simulated decays to a time (i.e., for decays).
In Sim 3, the Na source and the saber were positioned as in Run 3 and one billion decays were simulated. In the same way as for the thorium, we used the known activity of to convert the simulated number of decays to a time.
In Sim 4, we modeled the three sabers side-by-side, under the ADR, and produced ten billion secondary cosmic ray particles using the Cosmic Ray Library (CRY package [13]). These cosmic ray particles, as well as those particles produced by their interactions with the setup, hit the substrate and the sabers. The saber hits were disregarded if their energy deposition was lower than , the trigger threshold used during data acquisition.
From Sim 4, using the triple coincidence rate from Run 4A as a time conversion factor, we obtained a simulated cosmic ray DCR. This was then compared to the high-energy DCR observed in Run 1. We can see that cosmic rays constitute a major component of the high-energy DCR (see Fig. 6). Additionally, Sim 4 allowed us to calculate the ratios of quadruple-to-triple coincidences and quadruple-to-total high-energy TES hits. These ratios are analogous to those measured in Run 4A and Run 4B, respectively.
ICP-MS measurements. To investigate the contribution of radioactive decays to the number of high energy events seen by the TES we conducted an inductively coupled plasma-mass spectrometry (ICP-MS) analysis at the Laboratori Nazionali del Gran Sasso (LNGS, L’Aquila, Italy). This was to evaluate the intrinsic radioactivity of the materials around the TES, specifically the substrate and the copper box housing the TES. The analysis focused on uranium and thorium, which we suspected to be the main contributors of radioactive decays. In the substrate, uranium and thorium concentrations were found to be less than and , respectively. For the copper box, concentrations of both isotopes were below . Upon incorporating the measured radioactive activities of the copper box and substrate into our GEANT4 simulation, it was determined that the TES would register approximately one hit every 500 days (). This finding indicates that the impact of naturally occurring radioactivity on the high-energy DCR is negligible.

### v4

# Dark counts in optical superconducting transition-edge sensors for rare-event searches

@@ Abstract
I Introduction
Superconducting transition-edge sensors (TESs) are highly sensitive microcalorimeters, capable of detecting single- and multi-photon events with an efficiency close to unity [1]. Most TESs operate below , at the transition point between their superconducting and normal states, and can be tailored to different wavelength ranges, spanning from x-ray to infrared.
Several particle-physics experiments are employing—or plan to employ—TESs in rare event searches in the (0.1–10) eV range because of their low background and high detection efficiency. TESs sensitive to photons in this energy range are referred to as optical TESs. Examples of experiments using optical TESs include PTOLEMY [2, 3], targeting relic neutrinos, and ALPS II, focusing on new light bosons and other weakly interacting subelectronvolt particles [4]. Understanding the background of TESs is essential for this kind of rare-event searches. A previous study [5] investigated the background, also referred to as dark-count rate (DCR), for an infrared TES optimized for a wavelength of . While the authors identified various kinds of events contributing to the intrinsic DCR (counts detected by the TES without any input optics), they did not investigate the nature of these specific signals. They did, however, explore the DCR of a fiber-coupled TES, where the “warm” fiber end was located outside the cryostat (see Section 5.2 of Ref. [5]). It is important to note that this latter case represents the total dark-count rate of the device as defined by Ref. [6], which differs from the intrinsic DCR that is the focus of our current study.
In our work, we investigate the nature of the signals when no optical fiber is connected to the TES. Moreover, we measure the DCR over a broader wavelength range, spanning from 390 to . To assess the nature of the observed background events, we have devised distinct experimental layouts, employing radioactive sources and a cosmic ray coincidence setup.
@@ I Introduction
II Superconducting
transition-edge sensors
II.1 Design and operating principle
A TES is a thin superconducting film operated at the transition between its superconducting and normal state where it is most sensitive to changes in temperature. The TES working point is stabilized on the transition by negative electrothermal feedback [1] which is achieved by applying a voltage bias. This means that any increase in temperature due to energy absorption leads to an increase in resistance, which in turn causes a decrease in current. This drop in current reduces Joule heating, thereby cooling the sensor and bringing it back to the working point. A direct current superconducting quantum interference device (dc SQUID) connected to the TES acts as a transimpedance amplifier [10]. It converts the current drop resulting from the particle interaction to a measurable voltage drop. This voltage can then be amplified by standard electronic amplifiers and the waveforms can be measured and recorded by an oscilloscope. The voltage drop produced by the dc SQUID is proportional to the energy deposition in the TES within its operating range [1].
The two superconducting TESs used in this work were fabricated and operated at the Istituto Nazionale di Ricerca Metrologica (INRiM) in Turin (Italy). They are made of a bilayer of titanium, the superconducting material, and gold, a normal metal to lower the critical temperature, , of the titanium (approximately 400 mK) by the proximity effect [11]. Both TESs exhibit a of around and are operated inside an adiabatic demagnetization refrigerator (ADR). Since TESs and dc SQUIDs are sensitive to magnetic fields, including the geomagnetic field, a cryogenic magnetic shield is installed within the ADR around the working area to suppress any magnetic interference that could degrade the performance of the sensor. The magnetic shield is a cylindrical case made of a high-permeability nickel-iron alloy, manufactured by Amuneal, providing shielding capabilities [12].
Two 20 TESs were deposited on silicon substrates. We refer to the sensors as “TES E” and “TES B.” Most of the data were acquired from TES E, making it the primary focus of our study. Unless otherwise specified, references to TES data should be assumed to pertain to TES E. In Appendices A and D, we provide details on the TES fabrication and some information about the characterization and dark-count rate of TES B. For both TES E and TES B, the bath temperature was consistently maintained at , with the working point set at 15% of their normal resistance. The oscilloscope used for the data acquisition was a LeCroy Wave Runner. The TES signal has only been amplified by a single-stage 16-SQUID series array [13], using the XXF-1 Magnicon electronics [14].
II.2 Calibration
To calibrate the TES, we have coupled it to three different attenuated monochromatic lasers of 1540, 850, and using an optical fiber placed about one millimeter away from the sensor. The attenuation allows us to restrict the number of photons arriving at the TES and to avoid signal saturation, focusing on the region containing between one and eight photons. In Fig. 1, the calibration curve is shown on the right, where we plot the measured signal amplitude from the TES against the known energy of the impinging photon(s). The left side of the figure displays the histograms of counts versus amplitude for each laser wavelength used. We have taken the absolute minimum value as the pulse amplitude from the unfiltered waveforms. Thus, even in the absence of energy deposition in the TES, the minimum value is nonzero, as fluctuations in the baseline signal around zero occur. This results in a nonzero intercept for the calibration curve. We have intentionally not applied any filtering during the calibration process in order to match the unfiltered oscilloscope conditions during data acquisition.
III Event classification
Signals detected by the TES without any input optics are referred to as “intrinsic detector dark-counts” by Ref. [6]. However, in this paper we drop the adjective “intrinsic,” as we have identified external sources contributing to these counts such that not all of the dark-counts come from within the sensor (i.e., they are not intrinsic to it). The observed signals fall into three distinct categories: photonlike, high-energy, and electrical-noise events. While the three kinds of pulses are visually distinguishable, we have automated the categorization using the following pipeline developed in python. After applying a Butterworth filter, we calculate a set of variables for each waveform (e.g., full width at half maximum, maximum amplitude, pulse area, etc.). We employ principal component analysis (PCA) to find the two linear combinations of variables that maximize the variance of the points [15]. -means clustering is then applied to the resulting two-dimensional phase space to identify clusters of similar events [16]. To ensure the accuracy of this automated sorting, we have conducted a manual review of the clusters. This review process was facilitated by a graphical user interface (GUI) that allows for the selection of a cluster center, subsequently displaying all waveforms associated with that cluster for quick verification of the classification. In some cases, as demonstrated in Fig. 2, it was necessary to reapply the clustering algorithm on a subset of the data to fully separate the three groups.
The photonlike events are the green points in group A in Fig. 2, which were obtained by rerunning the classification algorithm on clusters 8 and 9. These signals are characterized by a sharp rise time and look identical to the ones we see when we shine light from the attenuated laser diode onto the TES.
The signals in clusters 1–7 in Fig. 2 are characterized by a long rise and decay time, suggesting that these events are caused by large energy depositions in the TES substrate, likely coming from secondary cosmic ray particles or naturally occurring radioactive decays. We therefore refer to them as “high-energy events.” This classification is consistent with the observed long rise time, which can be attributed to the slow diffusion of energy from the substrate into the TES.
The electrical-noise events (orange cluster B in Fig. 2 obtained by reanalyzing clusters 8 and 9) exhibit voltage spikes oscillating between positive and negative values and are characterized by a pulse area close to zero. We attribute these events to electromagnetic interference with the readout apparatus. The fact that they are still detectable even when the TES is in a fully superconducting state, a condition in which it is not sensitive to energy depositions, indicates that the culprit is not the TES.
IV Experimental Runs
The photonlike events are the expected signal for dielectric haloscopes searching for dark-matter dark photons. Therefore, it is essential to measure the photonlike DCR in the TES when the sensor is operated in the dark, with no optical link to ambient and in the absence of radioactive sources, as this forms the background against which the dielectric haloscope operates.
At the onset of our investigation, we expected to see a photonlike DCR on the order of Hz, in line with the findings by Ref. [5], the only preceding study of a similar nature. There was also a question of whether cosmic rays or natural radioactive decays, interacting with either the TES or its substrate, could account for some of these photonlike events as suggested by the same authors. To explore this possibility, and to gain a deeper understanding of the nature of high-energy events, which had not been extensively studied in optical TESs before, we have carried out four specific tests, explained below. It is important to note that all the experimental runs were conducted in the absence of a light source and without an optical fiber connected to the TES, which was entirely encased in a copper box thermally coupled to the 30-mK-stage (see Fig. 3). Although three fibers were present in the cryostat to allow for the initial calibration procedures, their ends were fixed to the outer side of the 3-K-stage within the ADR.
The first run was a background test (Run 1) intended to measure the photonlike DCR. Additionally, in the same run, we measured the high-energy DCR, which we hypothesized to be caused by cosmic rays or natural radioactive decays interacting with the TES substrate.
In Runs 2 and 3, we used two radioactive sources, 232Th and 22Na, respectively. These sources were placed against the exterior of the ADR and aligned perpendicularly to the TES substrate, which was mounted vertically inside the ADR (see Fig. 3). The aim of these tests was to examine the response of the TES to gamma events using radioactive sources. We expected that environmental gamma would yield a similar response in the sensor. In Run 2, a 232Th source with an activity measured at approximately was used. For Run 3, we employed a 22Na source, which had an activity of around . The emission spectrum of 22Na is characterized by two prominent photoelectric peaks. Notably, one of these peaks corresponds to the simultaneous emission of two back-to-back 511-keV photons, a phenomenon that occurs due to electron-positron annihilation following a decay. This source was sandwiched between the ADR and a plastic scintillator connected to a photomultiplier tube (PMT), which we call a “saber.” This setup allowed us to measure the double coincidences of the back-to-back gamma rays, capturing these occurrences with both the saber and the TES.
For the fourth test (Run 4), to demonstrate that cosmic rays can also induce high-energy events in the TES, we implemented a cosmic ray coincidence system, placing three sabers, side by side, directly below the ADR (see Fig. 4). A simultaneous triggering of the TES and all sabers indicated a cosmic ray shower or muon-bundle event. The three sabers were operated at the same gain within a 10% uncertainty. The threshold was set at on all three, which corresponds to (for more details about the calibration of the sabers, see Appendix F).
The data acquisition was conducted in two modes: in Run 4A, which lasted about 19 hours, the oscilloscope was configured to trigger on the AND signal of the three sabers, recording the TES signal simultaneously. The coincidence in Run 4A was set using the smart trigger function of the oscilloscope, requiring a signal to exceed the threshold in all the saber channels across the entire window. In Run 4B, spanning around 14 hours, the trigger was set on the TES for signals greater than , while recording the outputs of all three sabers. Run 4A has provided the ratio of quadruple coincidences (events registered by both the TES and the three sabers) to triple coincidences (events detected solely by the sabers), while Run 4B allowed us to estimate the ratio of quadruple coincidences to the total number of high-energy events detected by the TES.
V GEANT4 Simulation
Preliminary calculations indicate that direct cosmic ray interactions with the TES alone cannot account for the observed number of high-energy events. To estimate the maximum possible direct cosmic ray contribution, we have considered a scenario in which the TES is oriented horizontally, maximizing its exposure. At sea level, the average muon flux is about (see, for example, [17]). Even doubling this value to conservatively account for cosmic ray particles other than muons, a 20 TES would receive only around direct hits per day. In contrast, our larger 6.7, 500-µm-thick substrate would be impacted by around 400 cosmic ray particles per day. Since we observe about a thousand high-energy events per day in our measurements, far exceeding the estimated direct hits on the TES, we conclude that these events must (primarily) originate from energy depositions within substrate, and not direct interactions with the TES active area. Indeed, as we shall see later, these high-energy events arise from both cosmic rays and environmental gamma rays interacting with the substrate.
We have developed a geant4-based simulation to validate our understanding of the high-energy events seen by the sensor. We have ran the simulation in three configurations called Sims 2, 3, and 4, which mirror the experimental setups of Runs 2, 3, and 4, respectively. All simulations have used the substrate as the sensitive detector rather than the TES itself.
In Sim 2, we modeled the 232Th source as in Run 2 and simulated one billion decay events. The simulation assumed the source to be in secular equilibrium. This approach is supported by empirical data obtained from analyzing the thorium source with a high-purity germanium (HPGe) detector at NYUAD, which provided activity-level and spectral information consistent with this assumption. The activity from these measurements, around , was also used to convert the number of simulated decays to a time (i.e., approximately for decays).
In Sim 3, the 22Na source and the saber were positioned as in Run 3 and one billion decays were simulated. In the same way as for the thorium, we used the known activity of to convert the simulated number of decays to a time.
In Sim 4, we modeled the three sabers side by side, under the ADR, and generated ten billion secondary cosmic ray particles using the Cosmic-ray Shower Library (CRY package [18]). These cosmic ray particles, as well as those particles produced by their interactions with the setup, hit the substrate and the sabers. The saber hits were disregarded if their energy deposition was lower than , which corresponds to the trigger threshold used in Run 4A.
In addition to cosmic rays, we also simulated environmental gamma-rays originating from the radioactivity of building materials and other naturally occurring radiation sources. These gamma rays were generated according to the actual gamma-ray spectrum measured in the laboratory, using a NaI crystal coupled to a photomultiplier tube. Our measurements indicated an average gamma flux of gamma cm-2 s-1.
From Sim 4, we obtained a simulated high-energy DCR, which was then compared to the high-energy DCR observed in Run 1 (see the top green and orange points in Fig. 6). To accomplish this, we added together the rate of events detected in the TES substrate during the environmental gamma-ray simulation and the rate obtained during the cosmic ray simulation. To convert the simulated counts (the output from the simulation) to a simulated rate, we used the triple coincidence rate from Run 4A and the environmental gamma rate from the NaI detector as time-conversion factors. Additionally, Sim 4 enabled us to calculate the ratios of triple to quadruple coincidences and high-energy TES hits to quadruple coincidences. These ratios correspond to those measured in Runs 4A and 4B, respectively.
To investigate the contribution of radioactive decays to the number of high-energy events seen by the TES, we have conducted an inductively coupled plasma-mass spectrometry (ICP-MS) analysis at the Laboratori Nazionali del Gran Sasso (LNGS, L’Aquila, Italy). This was used to evaluate the intrinsic radioactivity of the materials around the TES, specifically the substrate and the copper box housing the TES. The analysis focused on uranium and thorium, which we suspected to be the main contributors of radioactive decays. In the substrate, the uranium and thorium concentrations were found to be less than (parts per billion) and , respectively. For the copper box, the concentrations of both isotopes were below . Upon incorporating the measured radioactivity of the copper box and substrate into our geant4 simulation, it was determined that the TES would register approximately one hit every 500 days (). This result indicates that the intrinsic radioactivity of the materials in close proximity to the TES contributes negligibly to its high-energy dark-count rate.
VI Results and discussion
[…]